OpenStudy (anonymous):

Calculate molality,molarity and mole fraction of KI if the density of 20%(mass) aqueous KI is 1.202 g/mL

OpenStudy (anonymous):

(a) Molar mass of KI = 39 + 127 = 166 g mol−1 20% (mass/mass) aqueous solution of KI means 20 g of KI is present in 100 g of solution. That is, 20 g of KI is present in (100 − 20) g of water = 80 g of water Therefore, molality of the solution = 1.506 m = 1.51 m (approximately) (b) It is given that the density of the solution = 1.202 g mL−1 ∴Volume of 100 g solution = 83.19 mL = 83.19 × 10−3 L Therefore, molarity of the solution = 1.45 M (c) Moles of KI Moles of water Therefore, mole fraction of KI = 0.0263

OpenStudy (anonymous):

@JOYAL u r a +2 student studying in kerala is it??